# Identifying and Addressing Delusions for Target-Directed Decision Making

### v4

@@ 4.2 Mixtures
future” Can be used to learn a conditional generator with temporal abstractions In addition to the shortcomings of “episode” (those for estimators only), this additionally causes E.3.2 when used as the exclusive strategy for estimator training Creates training data with temporally-ordered source-target pairs from the same episodes
@@
Table 1: Hindsight Relabeling Strategies & The Ideas Behind: “episode” and “future” are widely used as they increase sample efficiency in non-delusional cases significantly; “generate” and “pertask”, proposed in this paper, are effective against delusions, useful in specific scenarios.
@@ 4.3 Hybrid Strategies - A $2$-Slotted Approach
5 Related Works
Delusions in RL can stem from various other improper designs. For example, Lu et al. (2018) identified delusions caused by the limitations of function approximators for greedy policies in modelfree RL; Often accompanying hallucinations, delusions can arise in non-target-directed frameworks as well. In an offline planning framework such as Dyna, state hallucinations can irreversibly destabilize value estimations (Jafferjee et al., 2020). In Di Langosco et al. (2022), the authors classified goal misgeneralization, a class of delusional behaviors with which an agent competently pursues an undesired target leading in novel test situations. Zhao et al. (2024) gave first examples of delusional behaviors caused by improper hindsight relabeling. The perspective of delusions can often be overlooked because of the lack of estimators in target-directed methods, as discussions regarding hallucinations get magnified.
@@ 6 Experiments
To investigate the effectiveness of our strategies against delusions and how they can help target-directed agents achieve better generalization, we choose a combination of environments with clearly defined delusional cases, as well as methods whose estimators can be easily analyzed. This leads to our $4$ sets of experiments, coming from a combination of $2$ environments (one dominantly haunted by G.3.1, and another by G.3.1.1) and $2$ target-directed frameworks. Due to page limit, $3$ out of $4$ sets of experiments are only presented in the Appendix.
@@ 6.1 Environments & Settings
For environments, we favor tasks where the identified delusions are present and intuitive to inspect, for which the introduced SSM offers us clear advantage. Introduction and the results on another environment are presented in the Appendix.
For each training seed run, we sample and preserve $50$ training tasks from a distribution of tasks of size $12\times 12$ and difficulty $\delta=0.4$. All agents are trained for $1.5\times 10^{6}$ interactions by randomly selecting one of the $50$ tasks for each training episode. To speed up training, the initial state distributions span all the non-terminal states in each training task. Note that the change of initial states also increases risks of E.3.2.2, due to the presence of dense episode-terminating lava traps and relatively short MELs ($128$ for SSM).
@@ 6.2 Evaluative Criteria
We employ the following criteria to investigate the changes in agent’s estimations and behaviors:
Estimation Error - E.3.2, E.3.2.1, E.3.2.2 & Non-Delusional: At each evaluation point, we solve the ground truth distances between states and inspect the agents’ estimation errors. The errors are split based on the source-target pairs, into those involving problematic targets, i.e., E.3.2.1 and E.3.2.2, and those involving only valid targets, which includes the case of E.3.2.
Delusional Behavior Frequencies: We also monitor the frequency of a problematic target (G.3.1 or G.3.1.1) being chosen by the agents, as the result of their delusions in decision-time planning.
@@ 6.3 Methods
To clearly demonstrate the effectiveness against delusions and the resulting OOD performance boost, we adapted the following target-directed methods:
Skipper: At decision-time, it generates candidate target states that, together with the current state, form a directed graph. The states act as vertices and the edges are pairwise estimations of cumulative rewards and discounts (interchangeable with distances), under its evolving policy. A target state is chosen among candidates after applying value iteration (VI), i.e. the favorability of targets are defined by the $Q$ values of the resulting planned paths. Candidates can be preserved for re-selection of targets (Zhao et al., 2024).
LEAP (in Appendix): Instead of VI, LEAP uses an evolutionary algorithm to evolve sequences of subgoals leading to the task goal, with distances as fitness. The immediate subgoal of the elitest sequence is chosen to be the target, and used to condition a lower-level policy. LEAP is prone to delusions, since one problematic subgoal can destroy a whole sequence (Nasiriany et al., 2019).
For both agents, the generators and estimators are trained with HER, allowing easy switching of relabeling strategies for comparison. To a degree, both agents employ update rules that can lead to rejection of wrong estimations with data exposure (more discussions in Appendix). Targets are acquired by encoding imagined observations from end-to-end learned generators, where problematic propositions can be clearly identified (G.3.1 & G.3.1.1).
@@ 6.4 Hindsight Relabeling Strategies
Besides on the basic “episode”, “future” and “pertask”, we conduct comparative studies with the following baselines for fairer comparison (excluding bad generator performance):
@@
Additionally, the following hybrid strategies are proposed and compared (more in Appendix):
    F-(E+G): A hybrid strategy to deal with mostly E.3.2.1. “future” for generator, “episode” for estimator with $50\%$ chance using “generate” JIT, resulting in a mixture of $50\%$episode” and $50\%$generate” for estimator;
    F-(E+P): A hybrid strategy for mostly E.3.2.2. “future” for generator, mixture of $50\%$episode” and $50\%$pertask” for estimator;
    F-(E+P+G): A hybrid strategy against both E.3.2.1 and E.3.2.2 behaviors. “future” for generator, mixture of $2/3$episode” and $1/3$pertask” for estimator, with $1/4$ chance using “generate” JIT, resulting in a mixture of $50\%$episode”, $25\%$pertask” and $25\%$generate” for estimator;
@@
Figure 3: Skipper on SSM: each mean curve and CI (95%) bar are established over $20$ seed runs. Subfigures b,c,d,f,g,h) share the same legends. a-c) focus on E.3.2.1 delusions: a) Evolving ratio of G.3.1 targets among all candidates at each target selection, throughout training, is presented. Variants are differentiated only with the generator HER strategies; b) Evolving E.3.2.1 delusions measured by $L_{1}$ error in estimated distance throughout training is visualized. The distances to unreachable targets are clipped to be the maximum values that the estimator could output ($16$ in experiments). A more detailed breakdown of this error is in the Appendix; c) The curves represent the frequencies of choosing G.3.1 targets whenever a selection of targets is initiated; d) Final estimation accuracies of non-delusional source-target pairs after training completed, across a spectrum of ground truth distances. In this subfigure, both distances (estimation and ground truth) are conditioned on the snapshot of the final policies. Note that a part of the errors comes from E.3.2; Subfigures e-g) are the E.3.2.2-related counterparts of a-c); h) Each data point represents OOD evaluation performance aggregated over $4\times 20$ newly generated tasks, with mean difficulty matching the training tasks. The decomposed curves can be found in the Appendix.
6.5 Skipper on SSM (Set 1/4)
Generators: For the HER-trained generators, Fig. 3 a), shows that none among “future”, “episode” and “pertask” could directly address G.3.1 target generation with significance. While, Fig. 3 e) indicates that, “future” generates G.3.1.1 delusional targets significantly less frequently than “episode” and “pertask”, as the other two wasted training budget on G.3.1.1 targets, especially “pertask” that brings in even more problematic training samples from long distances. For fairer comparison in the estimators, we only compare variants with “future” for the generator training, i.e., “F-*”.
Estimators: Exclusive use of “episode” for estimators resulted in high accuracy in non-delusional cases (cases involving pairs of non-delusional states, “F-E” in Fig. 3 d)) for source-target pairs of both short and long distances, but low accuracies in E.3.2.1 and E.3.2.2 cases (Fig. 3 b) & f)); Exclusive use of “pertask” results in significantly worse non-delusional short-distance estimates (“F-P” in Fig. 3 d)), yet much better than “episode” (“F-E”) and “future” (“F-F”) in long-distance cases (indicating that trajectory-level “future” and “episode” lead to E.3.2 cases for longer-distance source-target pairs). This is likely because, without the presence of a backbone strategy such as “episode”, the estimator spent too much of its budget learning long-distance non-delusional or E.3.2.2 cases; Similarly, for “generate” (“F-G”), high accuracy is achieved for E.3.2.1 related cases, at the sacrifice of E.3.2.2 and non-delusional cases; All $3$ hybrid non-baseline strategies achieve good estimation accuracies in both short- and long-distance non-delusional estimation, as shown in Fig. 3 d). At the same time, we can observe particularly significant improvement in accuracy in E.3.2.2 delusional estimates in Fig. 3 f), by “F-(E+P)” and “F-(E+P+G)”, the hybrids assisted by “pertask”.
@@ 6.5 Skipper on SSM (Set 1/4)
Behaviors: For all compared variants, we can deduce from 3 that generally, less frequent G.3.1.1 generation (e)) and lower E.3.2.2 errors (f)) lead to less frequent delusional behaviors in g), which in turn improves the OOD performance in h); “episode” (“F-E”) showed mediocre evaluation performance (h)) for its good estimation accuracy in short-distance non-delusional cases, but was prone to delusional behaviors; While, despite that “pertask” (“F-P”) showed decent accuracies in longer-distance non-delusional and E.3.2.2 cases (d) & f)), its low accuracy in short-distance non-delusional estimation (d)) devastated the agent to exhibit the lowest performance in h); Similarly, despite that “generate”’s (“F-G”’s) effectiveness in addressing E.3.2.1-induced behaviors (Fig. 3 c)), its resulting bad estimation accuracies in the non-delusional cases destroyed its OOD performance.
All $3$ non-baseline hybrids achieve significantly better performance, as shown in Fig. 3 h). “F-(E+P)” and “F-(E+P+G)”, the hybrids assisted by “pertask”, performed the best in terms of reducing E.3.2.2 delusions (in Fig. 3 f)) and consequently addressed the most delusional behaviors (in Fig. 3 g)). This indicates a major presence of E.3.2.2 behaviors of Skipper’s failure modes on SSM.
@@ 6.6 Summary of Experiments
With the proposed strategies, we saw a reduction in delusional behaviors during training in both Skipper and LEAP (in Appendix), which led to better OOD generalization performance in two sets of environments, posing challenges of G.3.1 and G.3.1.1, respectively (more results in Appendix).
@@ 7 Empirical Guidelines
    Use update rules with appropriate properties and try to maximize the diversity of the training data distribution (against E.3.2);
@@
    Inspect the candidate targets if possible, since their low quality may indicate higher E.3.2.1 risks, which we can mitigate with the idea behind “generate”;
@@
    Analyze the state structure of the target tasks and identify temporary unreachabilities. These E.3.2.2 risks may indicate the effectiveness of the idea behind “pertask”.
@@ Appendix A Summaries
For better understanding of the contributions of this paper, we provide some brief summaries about the identified delusions / delusional behaviors and the corresponding mitigation strategies. For generator delusions, please check Tab. A. For estimator delusions, please check Tab. A. We hope that these could inspire the application on more target-directed methods.
@@
Table 2: Causes & Solutions for Generator Delusions
@@
Table 3: Causes & Solutions for Estimator Delusions
@@ B.2 Skipper on RDS (Set 2 / 4)
The second environment we employ is RandDistShift, abbreviated as RDS. RDS was originally proposed in Zhao et al. (2021) as a variant of the counterparts in the MiniGrid Baby-AI platform (Chevalier-Boisvert et al., 2018a; b; Hui et al., 2020) and then later used as the experimental backbone in Zhao et al. (2024). We can view RDS as a simpler version of SSM, where everything happens in semantic class $\langle 1,1\rangle$, i.e., agents always spawn with the sword and the shield in hand, thus can acquire the terminal sparse reward by simply navigating to the goal. RDS instances thus have smaller state spaces than its SSM counterparts. The most important difference, in the views of this paper, is that RDS removed the challenges introduced by temporary unreachabilities. This means that G.3.1.1 and E.3.2.2 are no longer relevant, shifting the dominance towards G.3.1 + E.3.2.1 combination. Using RDS not only showcase the performance of the proposed strategies on a controlled environment with G.3.1 + E.3.2.1 dominance, contrasting the G.3.1.1 + E.3.2.2 dominance of SSM, it also can be used to validate the performance of our adapted agents, on an environment where previous benchmarks exist.
@@ B.3 LEAP on SSM (Set 3 / 4)
For LEAP, we use some different metrics to analyze the effectiveness of the proposed strategies in addressing delusions. This is because, if LEAP’s estimator successfully addressed delusions and learned not to favor the problematic targets (G.3.1 and G.3.1.1), then they will not be selected in the evolved elitest sequence of subgoals. This makes it inconvenient for us to use the distance error in the delusional source-target pairs during decision-time as a metric to analyze the reduction of delusional estimates, because of their growing scarcity.
@@ B.5 Discussions & More Details of “generate” & “pertask
B.5.1 Implementation pf “pertask
pertask” takes the advantage for the fact that training is done on limited number of fixed task instances. We give each task a unique task identifier. At relabeling time, “pertask” samples observations among all the transitions marked with the same identifier as the current training task instance. This can be trivially implemented with individual auxiliary ERs that store only the experienced states with memory-efficient pointers to the buffered $x_{t}$’s in the main HER.

### v5

@@ 4.2 Mixtures
future” Can be used to learn a conditional generator with temporal abstractions In addition to the shortcomings of “episode” (those for estimators only), this additionally causes E.3.2 when used as the exclusive strategy for estimator training Creates training data with temporally ordered source-target pairs from the same episodes
@@
Table 1: Hindsight Relabeling Strategies & The Ideas Behind: episode” and “future” are widely used as they increase sample efficiency in non-delusional cases significantly; “generate” and “pertask”, proposed in this paper, are effective against delusions, useful in specific scenarios.
@@ 5 Experiments
To investigate the effectiveness of our strategies against delusions and how they help target-directed agents achieve better generalization, we implement a combination of environments with clearly defined delusional cases, as well as methods whose estimators can be easily analyzed. This leads to our $4$ sets of experiments, coming from a combination of $2$ environments (one dominantly haunted by G.3.1, and another by G.3.1.1) and $2$ target-directed frameworks. Due to page limit, $3$ out of $4$ sets of experiments are presented in the Appendix. We summarize those at the end of this section.
@@ 5.1 Environments & Settings
For environments, we favor tasks where the delusions are present, dangerous and intuitive to inspect, for which the introduced SSM offers us clear advantage. Introduction and the results on another environment are presented in the Appendix.
For each training seed, we sample and preserve $50$ training tasks of size $12\times 12$ and difficulty $\delta=0.4$. All agents are trained for $1.5\times 10^{6}$ interactions by randomly selecting one of the $50$ frozen tasks for each training episode. To speed up training, we make the initial state distributions span all the non-terminal states in each training task. This change increases risks of E.3.2.2, due to the presence of dense episode-terminating lava traps and relatively short MELs ($128$ for SSM).
@@ 5.2 Evaluative Criteria
We use the following criteria to investigate the changes in agent’s estimations and behaviors:
Estimation Error - E.3.2, E.3.2.1, E.3.2.2 & Non-Delusional: At each evaluation point, we solve the ground truth distances between states and inspect the agents’ estimation errors. The errors are split based on the source-target pairs, into those involving problematic targets, i.e., E.3.2.1 and E.3.2.2, and those involving only valid targets, which includes the case of E.3.2. The estimation errors show the internal degree of delusions about targets within the estimators.
Delusional Behavior Frequencies: We also monitor the frequency of a problematic target (G.3.1 or G.3.1.1) being chosen by the agents, as the result of their delusions in decision-time planning. The frequencies demonstrate how often delusional targets are being favored because of delusions.
@@ 5.3 Methods
We seek to demonstrate the generality of our strategies by applying them onto different methods with very different behaviors. Thus, we adapted the following target-directed methods:
Skipper: generates candidate target states that, together with the current state, form a directed graph at decision-time. The states act as vertices and the edges are pairwise estimations of cumulative rewards and discounts (interchangeable with distances), under its evolving policy. A target state is chosen after applying value iteration (VI), i.e. the favorability of targets are defined by the $Q$ values of the planned paths (Zhao et al., 2024).
LEAP (in Appendix): Instead of VI, LEAP uses an evolutionary algorithm to evolve sequences of subgoals leading to the task goal, with distances as fitness. The immediate subgoal of the elitist sequence is chosen to be the target, and used to condition a lower-level policy. LEAP is prone to delusions, since one problematic subgoal can destroy a whole sequence (Nasiriany et al., 2019).
For both agents, the generators and estimators are trained with HER, allowing easy switching of relabeling strategies for comparison. Roughly, both agents employ update rules that can reject wrong estimations with data exposure (more discussions in Appendix). Targets are acquired by encoding observations imagined by the generators, where G.3.1 & G.3.1.1 propositions are clearly identified.
@@ 5.4 Hindsight Relabeling Strategies
We focus on the following baselines for fairer comparison (excluding bad generator performance):
@@
Additionally, the following hybrid strategies are proposed and compared:
    F-(E+G): A hybrid strategy mostly for E.3.2.1. “future” for generator, “episode” for estimator with $50\%$ chance using “generate” JIT, resulting in a half-half mixture of “episode” & “generate”;
    F-(E+P): mostly for E.3.2.2. “future” for generator, half “episode” & half “pertask” for estimator;
    F-(E+P+G): for both E.3.2.1 & E.3.2.2 behaviors. “future” for generator, mixture of $2/3$episode” and $1/3$pertask” for estimator, with $1/4$ chance using “generate” JIT, resulting in a mixture of $50\%$episode”, $25\%$pertask” and $25\%$generate” for estimator;
5.5 Skipper on SSM (Set 1/4)
Generators: We first verify the generator’s contributions to delusional behaviors. For the HER-trained generators, Fig. 3 a), shows that none among “future”, “episode” and “pertask” could significantly address G.3.1 target generation, validating our claims about the challenges of hallucination. While, Fig. 3 e) indicates that, “future” generates G.3.1.1 delusional targets significantly less frequently than “episode” and “pertask”, as the other two wasted training budget on G.3.1.1 targets, especially “pertask” that brings in more problematic training samples from long distances. For fairer comparison in the estimators, we only compare variants with “future” for the generator training, i.e., “F-*”.
Estimators: Now, we look into the degrees of delusions inside the learned estimators. Exclusive use of “episode” for estimators resulted in high accuracy in non-delusional cases (cases involving pairs of non-delusional states, “F-E” in Fig. 3 d)) for source-target pairs of both short and long distances, but low accuracies in E.3.2.1 and E.3.2.2 cases (Fig. 3 b) & f)). This shows the source of popularity of these baseline strategies in existing literature; Unsurprisingly, exclusive use of “pertask” results in significantly worse non-delusional short-distance estimates (“F-P” in d)), yet much better than “episode” (“F-E”) and “future” (“F-F”) in long-distance cases (indicating that trajectory-level “future” and “episode” lead to E.3.2 cases for longer-distance source-target pairs). This is likely because, without the presence of a backbone strategy such as “episode”, the estimator wastes its budget learning long-distance non-delusional or E.3.2.2 cases; Demonstrating a similar tradeoff, for “generate” (“F-G”), high accuracy is achieved for E.3.2.1 related cases, at the sacrifice of E.3.2.2 and non-delusional cases; All $3$ hybrid non-baseline strategies achieve good estimation accuracies in both short- and long-distance non-delusional estimation, as shown in d). At the same time, we can observe particularly significant improvement in accuracy in E.3.2.2 delusional estimates in f), by “F-(E+P)” and “F-(E+P+G)”, the hybrids assisted by “pertask”. These indicate their effectiveness.
@@ 5.5 Skipper on SSM (Set 1/4)
Figure 3: Skipper on SSM: each mean curve and CI (95% for all subfigures except c) & g), which used 50% due to the chaotic overlap) bar are established over $20$ seed runs. Legends are shared among b,c,d,f,g,h). a-c) focus on E.3.2.1 delusions: a) Evolving ratio of G.3.1 targets among all candidates at each target selection, throughout training. Variants are differentiated only with the generator HER strategies; b) Evolving E.3.2.1 delusions measured by $L_{1}$ error in estimated distance throughout training. The distances to unreachable targets are clipped to be the maximum values that the estimator could output ($16$ in experiments). A more detailed breakdown of this error is in the Appendix; c) The curves represent the frequencies of choosing G.3.1 targets whenever a selection of targets is initiated; d) Final estimation accuracies of non-delusional source-target pairs after training completed, across a spectrum of ground truth distances. Both estimated and true distances are conditioned on the policies of the last timestep. Note that a part of the errors comes from E.3.2; Subfigures e-g) are the E.3.2.2-counterparts of a-c); h) Each data point represents OOD evaluation performance aggregated over $4\times 20$ newly generated tasks, with mean difficulty matching training. The decomposed curves are presented in the Appendix.
@@
Behaviors: Finally, we examine how reducing delusions affect the agents’ behaviors. For all compared variants, we can deduce from Fig. 3 that generally, less frequent G.3.1.1 generation (e)) and lower E.3.2.2 errors (f)) lead to less frequent delusional behaviors in g), which in turn improves the OOD performance in h); “episode” (“F-E”) showed mediocre OOD performance (h)) for its good estimation accuracy in short-distance non-delusional cases, but was prone to delusional behaviors; While, despite that “pertask” (“F-P”) showed decent accuracies in longer-distance non-delusional and E.3.2.2 cases (d) & f)), its low accuracy in short-distance non-delusional estimation (d)) devastated the baseline to the lowest performance in h); Similarly, despite “generate”’s (“F-G”’s) effectiveness in addressing E.3.2.1-behaviors (c)), its resulting bad estimation accuracies in the non-delusional cases destroyed its OOD performance. In contrast, all $3$ hybrids achieve significantly better OOD performance in h). Assisted by “pertask”, “F-(E+P)” and “F-(E+P+G)” performed the best in reducing E.3.2.2 delusions (in f)) and consequently addressed the most delusional behaviors (in g)). This indicates a major presence of E.3.2.2 behaviors of Skipper’s failure modes on SSM.
@@ 5.6 Summary of Experiments
With the proposed strategies, we saw a reduction in delusions in terms of estimation errors and in delusional behaviors in both Skipper and LEAP (in Appendix, more prone to delusions), which led to better OOD generalization performance in $2$ sets of environments, posing challenges of G.3.1 & G.3.1.1, respectively (more results in Appendix). All $4$ sets of experiments align in terms of conclusions.
6 Related Works
Delusions in RL can stem from other improper designs. Lu et al. (2018) identified delusions caused by the limitations of function approximators for greedy policies in model-free RL; Often accompanying hallucinations, delusions can arise in non-target-directed frameworks as well. In an offline planning framework such as Dyna, hallucinations can irreversibly destabilize value estimations (Jafferjee et al., 2020). Di Langosco et al. (2022) classified goal misgeneralization, a class of delusional behaviors with which an agent competently pursues an undesired target leading in novel test situations. Zhao et al. (2024) gave first examples of delusional behaviors caused by improper hindsight relabeling. The perspective of delusions can often be overlooked because of the lack of estimators in target-directed methods, as discussions regarding hallucinations get magnified.
@@ 7 Empirical Guidelines
    Use proper update rules and try to maximize the diversity of the training data (against E.3.2);
@@
    Inspect the candidate targets if possible, as their low quality may indicate higher E.3.2.1 risks, which can be mitigated with “generate”;
@@
    Analyze the state structure of the target tasks and identify temporary unreachabilities. These E.3.2.2 risks may indicate the effectiveness of “pertask”.
@@ Appendix A Summaries
For better understanding of the contributions of this paper, we provide some brief summaries about the identified delusions / delusional behaviors and the corresponding mitigation strategies. For generator delusions, please check Tab. A. For estimator delusions, please check Tab. A. We hope that these could inspire the application of more target-directed methods.
@@
Table 2: Causes & Solutions for Generator’s Contribution to Delusions
@@
Table 3: Causes & Solutions for Estimator’s Contribution to Delusions
@@ B.2 Skipper on RDS (Set 2 / 4)
The second environment we employ is RandDistShift, abbreviated as RDS. RDS was originally proposed in Zhao et al. (2021) as a variant of the counterparts in the MiniGrid Baby-AI platform (Chevalier-Boisvert et al., 2018a; b; Hui et al., 2020), and then later used as the experimental backbone in Zhao et al. (2024). We can view RDS as a simpler version of SSM, where everything happens in semantic class $\langle 1,1\rangle$, i.e., agents always spawn with the sword and the shield in hand, thus can acquire the terminal sparse reward by simply navigating to the goal. RDS instances thus have smaller state spaces than its SSM counterparts. The most important difference, in the views of this paper, is that RDS removed the challenges introduced by temporary unreachabilities. This means that G.3.1.1 and E.3.2.2 are no longer relevant, shifting the dominance towards G.3.1 + E.3.2.1 combination. Using RDS not only showcase the performance of the proposed strategies on a controlled environment with G.3.1 + E.3.2.1 dominance, contrasting the G.3.1.1 + E.3.2.2 dominance of SSM, it also can be used to validate the performance of our adapted agents, on an environment where previous benchmarks exist.
@@ B.3 LEAP on SSM (Set 3 / 4)
For LEAP, we use some different metrics to analyze the effectiveness of the proposed strategies in addressing delusions. This is because, if LEAP’s estimator successfully addressed delusions and learned not to favor the problematic targets (G.3.1 and G.3.1.1), then they will not be selected in the evolved elitist sequence of subgoals. This makes it inconvenient for us to use the distance error in the delusional source-target pairs during decision-time as a metric to analyze the reduction of delusional estimates, because of their growing scarcity.
@@ B.5 Discussions & More Details of “generate” & “pertask
B.5.1 Implementation of “pertask
pertask” takes the advantage of the fact that training is done on limited number of fixed task instances. We give each task a unique task identifier. At relabeling time, “pertask” samples observations among all the transitions marked with the same identifier as the current training task instance. This can be trivially implemented with individual auxiliary experience replays that store only the experienced states with memory-efficient pointers to the buffered $x_{t}$’s in the main HER.